OpenStudy (anonymous):

find the critical numbers of 4y^3 +2y + 6

OpenStudy (anonymous):

In the second term is it y or y^2

OpenStudy (anonymous):

If it is y itself then no critical pints since first derivative is never zero in the set of real numbers

OpenStudy (anonymous):

original function is (-y^2)^2 +y^2+6y+9 f '(y) = 4y^3+2y+6 f ''(y) = 12y^2+2

OpenStudy (anonymous):

Ohhh u have to mention than first dude I thought that it is the main function

OpenStudy (anonymous):

One of the critical point is -1

OpenStudy (anonymous):

i figured that after i read your response.. my bad..

OpenStudy (anonymous):

Its k

OpenStudy (anonymous):

im just having issues working out the first derivative to find my critical numbers..

OpenStudy (anonymous):

And the other two are imaginary

OpenStudy (anonymous):

differentiate the equation first 12y^2+2 then equate it to 0 12y^2+2=0 y^2= -1/6 aww man!!!!!!!!!!!!!!!!! its critical points doesn't exits

OpenStudy (anonymous):

DHASHINI once read our replies tooo u will get the info about the question

OpenStudy (anonymous):

ok.. my math could be wrong then. perhaps i should had started with my problem. find the point on the parabola x + y^2 =0 that is closest to (0,-3)

OpenStudy (anonymous):

Dude what you did was absolutely correct just go on we have only one critical point this may be the shortest too wait a second i will let u know

OpenStudy (anonymous):

Ya it is the point of minimum because the value of y=-1 gives positive for second derivative and substitute this value of y dude

OpenStudy (anonymous):

Point is (-1,-1) And surely it lies on the parabola also dude

OpenStudy (anonymous):

and the shortest distance is 5 units

OpenStudy (anonymous):

ok.. thanks.. i was certain the word point was a typo.. i was counting on the parabola to have 2 point(s)

OpenStudy (anonymous):

no probs